OpenStudy (anonymous):

The sum of the x-coordinates of the vertices of a triangle in the Cartesian plane equals 10. Find the sum of the x-coordinates of the midpoints of the sides of the triangle.

OpenStudy (anonymous):

10

OpenStudy (mathmate):

(1/2)(xa+xb)+(1/2)(xb+xc)+(1/2)(xc+xa) =(1/2)(2xa+2xb+2xc) =(xa+xb+xc) =10